Earley Corporation issued perpetual preferred stock with a 9% annual dividend. The stock currently yields 7%, and its par value is $100. Round your answers to the nearest cent.

a. What is the stock's value?

b. Suppose interest rates rise and pull the preferred stock's yield up to 14%. what is the new market value?
Answer #1

Annual dividend=100*9%=9

a.Stock value=Annual dividend/Current yield

=9/0.07=$128.57(Approx).

b.Stock value=9/0.14=$64.29(Approx).